Operating Systems and Systems Integration

Installing the Acrobat Reader and Browser Plugin

1 Aim

This is a quick guide to installing the Adobe\begin{matrix}c \\ ◯ \end{matrix} Acrobat reader and plugin that is on the powertools cdrom for Red Hat Linux.

2 Procedure

There are two steps here:

• Mount the network drive

• Install the Acrobat rpms

Here is how we do it. The procedure is very similar to installing updates, so refer to the document Applying OS Updates for more information. Note that here we are using sudo. Please read the work sheet about sudo first.

1. Type:

     $ sudo mkdir /mnt/ftp
     $ sudo mount nicku.org:/var/ftp/pub /mnt/ftp

Now you have mounted the directory /var/ftp/pub from the machine ictlab over the network to the local directory /mnt/ftp, using the nfs protocol. If you change to your local directory /mnt/ftp, you will be accessing the directory

          var/ftp/pub on ictlab over the network.

You can see all the directories exported from Ictlab by typing

          $ showmount -e ictlab

2. Now change to a directory on that network drive:

     $ cd /mnt/ftp/redhat-7.0/powertools

3. Close Netscape if it is running so that the plugin is installed correctly.

  4. Install the two RPM packages, acroread-4.05-6.i386.rpm and acroread-plugin-
     4.05-6.i386.rpm:
     $ sudo rpm -Uhv acroread*.rpm

Installing the Acrobat Reader and Browser Plugin Operating Systems and Systems Integration

2

The command rpm is important to you. There is a whole book about this command, called Maximum RPM, freely available on the documentation cdrom as the package max-rpm-en-1.0.noarch.rpm. You can get a copy here at

http://nicku.org/ftp/roswell-beta-2/docs/RedHat/RPMS/maximum-rpm-1.0-0.20010810.noarch.rpm.

The three options that we used here with the rpm program:

-U Update the package. If an older version ofthe package is already installed, uninstall the older version first, then install the new one. If an older version of the package is not installed, just install this new package.

-h Print hashes (#) to indicate progress. -v verbosely print the name of each package as install it.

That’s all. Open Netscape and click on a pdf document. If your screen resolution is set low, it may be difficult to click on the Accept button on Adobe’s eula. You may be able to reach it in the workspace below.